OpenStudy (anonymous):

log4 x=1/2 help please

OpenStudy (anonymous):

log(x+2) = 1 + log(x-1), assume log is base 4 x + 2 = 4^(1 + log(x-1)) = 4^1 * (x-1) = 4x - 4 x + 2 = 4x - 4 3x = 6 x = 2